# Do Vision Encoders Truly Explain Object Hallucination?: Mitigating Object Hallucination via Simple Fine-Grained CLIPScore

### v3

@@ Abstract
Recently, Large Vision-Language Models (LVLMs) show remarkable performance across various domains. However, these models suffer from object hallucination. This study revisits the previous claim that the cause of such hallucinations lies in the limited representational capacity of the vision encoder. Our analysis implies that the capacity of the vision encoder is not necessarily a major limiting factor in detecting object hallucination. Based on this insight, we propose Fine-grained CLIPScore (F-CLIPScore), a simple yet effective evaluation metric that enhances object-level granularity by incorporating text embeddings at the noun level. Evaluations on the OHD-Caps benchmark show that F-CLIPScore significantly outperforms conventional CLIPScore in accuracy by a large margin of 39.6% without additional training. We further demonstrate that F-CLIPScore-based data filtering reduces object hallucination in LVLM (4.9% in POPE).
@@ 1 Introduction
Recent studies identify Large Vision-Language Models (LVLMs) as a leading approach for vision-language integration (Liu et al., 2023; Wang et al., 2024a; Zhu et al., 2023; Chen et al., 2024). However, like hallucinations in large language models (LLMs) (Ji et al., 2023; Zhang et al., 2023), LVLMs exhibit object hallucination, referring to nonexistent or misidentified objects that may undermine their reliability (Li et al., 2023a; Liu et al., 2024b).
Liu et al. (2024c) built OHD-Caps, a dataset designed to measure object hallucination. The dataset comprises 1.5k image and captions pairs where a model needs to select the best caption that does not show hallucinations. They found that CLIPScore (Hessel et al., 2021) achieved only 10–20% accuracy, and the further fine-tuning with the proposed objective function with the dataset improves the accuracy up to 80–90%. However, when they connected the OHD-Caps-trained CLIP to an LVLM and conducted full fine-tuning, the resulting accuracy sometimes drops, showing lower performance compared to original CLIP (for instance, 1st row of Table 4 in (Liu et al., 2024c) shows the accuracy in POPE benchmarks (Li et al., 2023b) drops from 85.4% to 81.2%).
@@
To address this issue, we introduce Fine-grained CLIPScore (F-CLIPScore), a novel image-text correlation metric. F-CLIPScore leverages a sentence parser like spaCy Honnibal et al. (2020) and the forward pass of a Vision-Language Model (VLM) like CLIP Radford et al. (2021), offering an efficient way to evaluate the vision encoder’s representational capacity. Our experimental results show that applying F-CLIPScore to the OHD-Caps test set improves accuracy by +39.6% without additional training. This indicates that the limited capacity of the vision encoder may not be the primary cause of object hallucination. Additionally, we verify that using F-CLIPScore for pretraining data curation in LVLMs enables the training of models with reduced hallucination, even with significantly fewer data samples. Notably, in LVLM pretraining, data filtering alone improved POPE accuracy by 4.9% compared to the baseline.
@@
  • We provide evidence that object hallucination does not primarily stem from the vision encoder’s capacity.
@@ 2 Related Work
Object hallucination refers to cases in which the generated textual descriptions include objects that do not correspond to the given image (Liu et al., 2024b). LVLMs generally consist of three components: a vision encoder, an LLM, and an adapter (Liu et al., 2023). The structural characteristics of LVLMs contribute to object hallucination, which arises from multiple intertwined factors (Liu et al., 2024b). While some studies argue that hallucinations can be mitigated by enhancing the decoding process of the LLM (Manevich & Tsarfaty, 2024; Wang et al., 2024b), others suggest that one of the causes lies in the limited representational capacity of the vision encoder (Liu et al., 2024c). Additionally, some research indicates that training the adapter with contrastive data is essential to reduce object hallucination (Jiang et al., 2024). Moreover, the trained bias of the model has also been identified as a cause of hallucination (Hu et al., 2023; Liu et al., 2024a).
CLIPScore (Hessel et al., 2021) is a reference-free evaluation metric that assesses the consistency between an image and text caption by computing the cosine similarity between the embeddings generated by the vision encoder and text encoder of the CLIP model. Beyond its application in measuring caption quality, several studies have also leveraged CLIPScore for data curation in the training of Vision-Language Models (VLMs) (Schuhmann et al., 2021; Gadre et al., 2023).
A recent study utilized CLIPScore to evaluate object hallucination in Vision-Language Models (VLMs) (Liu et al., 2024c). Their findings suggest that this phenomenon stems from the limited capacity of the vision encoder. In this study, we carefully reassess this claim and demonstrate that object hallucination is not necessarily caused by the limitations of the vision encoder alone.
@@ 3.1 Motivation
We first computed the distribution of cosine similarity between the image embeddings from CLIP-L and OHD-Caps-trained CLIP-L, focusing on the samples that were correctly predicted by CLIP-L. We then performed the same analysis on the samples that were incorrectly predicted by CLIP-L and compared the two distributions. As shown in Figure 2a, we observe that there is no substantial difference between the two distributions. We conducted the same analysis using ground-truth text embeddings of the pre-trained and fine-tuned models, and similarly found no significant difference in the cosine similarity distributions between the samples that were correctly and incorrectly predicted by CLIP-L (Figure 2b). In contrast, Figure 2c reveals significant changes in text embeddings for captions with hallucination (green), and without hallucination (purple) highlighting the distinct adaptation of text representation. These observations suggest that OHD-Caps training induces more noticeable changes in the text representation space, particularly in discriminating hallucinated from non-hallucinated captions, whereas the vision representation space exhibits relatively minor changes.
@@ 3.2 Fine-grained CLIPScore
Motivated by these observations, we propose a simple metric called Fine-grained CLIPScore (F-CLIPScore), which enhances the discriminative power of the VLM in order to utilize textual information with more granularity without additional training. F-CLIPScore first utilizes the spaCy parser Honnibal et al. (2020) to extract nouns from a given sentence. Then, it evaluates the quality of an image caption by averaging the CLIPScore of the entire sentence and each individual noun (Figure 3). Mathematically, given a full sentence $S$ and a total of $N$ nouns, denoted as $n_{i}$, F-CLIPScore is defined as
$\small\operatorname{F-CLIPScore}(S)=\frac{\operatorname{CLIPScore}(S)+\sum_{i=1}^{N}{\operatorname{CLIPScore}(n_{i})}}{N+1}$ (1)
@@ 4 Experiments
For evaluating OHD-Caps test set, F-CLIPScore only requires an additional parsing step (we used en_core_web_sm of spaCy (Honnibal et al., 2020)), which takes on average 170 ms per sentence and can be parallelized across samples. We note that this parser runs in linear time with respect to the length of the sentence (Honnibal & Johnson, 2015). For training CLIP with equation 2 on the OHD-Caps train set, we used a batch size of 64 and a learning rate of 1e-5, which required 3 hours on an H100 GPU. For LLaVA pretraining, we employed an effective batch size of 128 and a learning rate of 1e-3, which took 7 hours on two H100 GPUs.
@@ 5.1 F-CLIPScore on OHD-Caps
We evaluated the OHD-Caps test set, an object hallucination assessment dataset, using the proposed F-CLIPScore. As shown in Table 1, evaluation results with OpenAI CLIP ViT-L Radford et al. (2021) indicate that F-CLIPScore outperformed the baseline model by up to 39.6% without additional training (row 4 vs. 5). However, it still performed 17.6% to 37.4% worse than trained models (row 5 vs. 7). This may indicate that although CLIP vision encoders may not be the main cause of object hallucination, further training may enhance their capability. Further validation of F-CLIPScore can be found in the Appendix.
@@
$\dagger$: We trained openai/clip-vit-large-patch14 with the train code from Liu et al. (2024c)
@@
To test whether F-CLIPScore is orthogonal to the previously proposed method Liu et al. (2024c), we modify the loss as
@@
where $L_{CLIP}$ is the contrastive loss proposed in  Radford et al. (2021), and $L_{OHD}$ is the marginal loss proposed in  Liu et al. (2024c). $B$ denotes the batch size, while $I_{i}$ and $C_{i}$ are image and caption embeddings of the $i$-th positive pair. $\alpha$ is a hyperparameter that we set to $0.3$. Applying F-CLIPScore as a loss improved performance by 0.7 percentage points on COCO and 0.1 percentage points on NoCaps (row 7 vs. 9).
5.2 LVLM Pretrain Data Curation with F-CLIPScore
As shown in Section 5.1, F-CLIPScore was able to exhibit competent performance in detecting object hallucination without any training, and could further serve as an orthogonal method for training VLMs. We aimed to investigate whether F-CLIPScore can influence the pretraining process of LVLM by acting as a data filter between the vision encoder and the LLM. To this end, we applied F-CLIPScore to filter the pretraining data for LLaVA (Liu et al., 2023), which consists of 558k samples, by removing the bottom $x\%$ of the alignment training set based on F-CLIPScore.
POPE Acc ($\uparrow$, %) Filtering rate (%)
Filtering Method 20 30 40 50 60
Random 50.7 49.9 49.9 50.5 49.8
CLIPScore (w/o train) 52.2 47.3 50.0 53.1 50.0
F-CLIPScore (w/o train) 51.6 55.5 50.6 50.1 51.8
CLIPScore (trained) 49.8 49.8 50.8 49.8 52.6
w/o filtering 50.6
Table 2: Accuracy on the POPE benchmark after LLaVA pretraining (vision-text alignment training before SFT). We use CLIP-L Radford et al. (2021) as a vision encoder, and Llama 2 7B Touvron et al. (2023) as an LLM backbone. “trained” indicates the OHD-Caps-trained CLIP-L. “Random” refers to the removal of $x\%$ of samples chosen at random.
As shown in Table 2, training the alignment model on the top 70% of data curated by F-CLIPScore resulted in a +4.9% improvement in POPE accuracy compared to training on the entire dataset (row 5 vs. 7). In contrast, using the OHD-Caps-trained CLIP for filtering yielded only marginal gains, and random filtering showed no improvement. These results may suggest that F-CLIPScore effectively captures object hallucination-related quality, even in general-purpose datasets. This finding underscores the need to explore alternative causes of object hallucination beyond the capacity of the vision encoder. Although model performance shows a non-linear trend with respect to the filtering ratio, our experiments used a fixed number of training epochs, which may account for this behavior. As Goyal et al. (2024) highlights the need to balance data quality with computational cost, identifying the optimal filtering ratio under such trade-offs remains an important direction for future research.
6 Conclusion
Limitations
Second, our method faces linguistic constraints and challenges in multilingual generalization. This study employs the spaCy parser (Honnibal et al., 2020) to extract nouns from text, a technique that performs relatively reliably in well-structured languages such as English. However, parsing accuracy may vary across different languages, potentially leading to inconsistencies in F-CLIPScore computation. To address this, future research should explore the scalability of F-CLIPScore by evaluating its effectiveness on multilingual datasets and refining the parsing methodology for broader linguistic applicability.
Appendix A Additional Results on F-CLIPScore Robustness
We randomly replaced the caption embeddings into noun embeddings extracted from a Wikipedia corpus (Davies, 2015) to examine how F-CLIPScore responds to such perturbations. As shown in Table 3, we observed the expected decline in accuracy as nouns were replaced. This indicates that the strong performance of F-CLIPScore, even without additional training, is not due to random noise, but rather stems from its ability to more effectively leverage fine-grained textual embeddings.
@@ Appendix A Additional Results on F-CLIPScore Robustness
Table 4: Accuracy on the OHD-Caps test set evaluated with OpenAI CLIP-L. $\operatorname{F-CLIPScore}_{N}$ refers to the noun-based $\operatorname{F-CLIPScore}$ defined in Equation 1. $\operatorname{F-CLIPScore_{NP}}$ uses noun phrases instead of individual nouns, and $\operatorname{F-CLIPScore}_{V}$ replaces the nouns with verbs.
@@
CECLIP (Zhang et al., 2024) 32.8 28.0 25.0
NegCLIP (Yuksekgonul et al., 2023) 52.8 40.8 23.4
@@
Our F-CLIPScore even outperforms the training-based methods (Zhang et al., 2024; Yuksekgonul et al., 2023), as shown in Table 5.
@@
SigLip ViT-L (CLIPScore) (Zhai et al., 2023) 48.6 38.4 30.2
@@
Licenses for the datasets and models used

### v4

@@ Abstract
Recently, Large Vision-Language Models (LVLMs) show remarkable performance across various domains. However, these models suffer from object hallucination. In this work, we study object hallucination primarily in a discriminative, retrieval-style evaluation setting (OHD-Caps), rather than in free-form caption generation. This study revisits the previous claim that the cause of such hallucinations lies in the limited representational capacity of the vision encoder. Our analysis implies that the capacity of the vision encoder is not necessarily a major limiting factor in detecting object hallucination. Based on this insight, we propose Fine-grained CLIPScore (F-CLIPScore), a simple yet effective evaluation metric that enhances object-level granularity by incorporating text embeddings at the noun level. Evaluations on the OHD-Caps benchmark show that F-CLIPScore significantly outperforms conventional CLIPScore in accuracy by a large margin of 39.6% without additional training. We further demonstrate that F-CLIPScore-based data filtering reduces object hallucination in LVLM (4.9% in POPE accuracy after alignment pretraining). Our code is publicly available at https://github.com/abzb1/f-clip
@@ 1 Introduction
Recent studies identify Large Vision-Language Models (LVLMs) as a leading approach for vision-language integration (LLaVA; qwen2vl; minigpt4; internvl). However, like hallucinations in large language models (LLMs) (llm_hallucination; sirenssong), LVLMs exhibit object hallucination, referring to nonexistent or misidentified objects that may undermine their reliability (eval_obj_hallucination_lvlm; lvlm_hallucination_survey). While object hallucination is often discussed in the context of caption generation, this paper focuses on discriminative vision–language models and retrieval-style evaluation protocols, where the goal is to select or score candidate captions given an image.
ohd-caps built OHD-Caps, a dataset designed to measure object hallucination in a discriminative caption-selection setting. The dataset comprises 1.5k image-captions pairs where a model needs to select the best caption that does not show hallucinations. They found that CLIPScore (clipscore) achieved only 10–20% accuracy, and the further fine-tuning with the proposed objective function with the dataset improves the accuracy up to 80–90%. However, when they connected the OHD-Caps-trained CLIP to an LVLM and conducted full fine-tuning, the resulting accuracy sometimes drops, showing lower performance compared to original CLIP (for instance, 1st row of Table 4 in (ohd-caps) shows the accuracy in POPE benchmarks (pope) drops from 85.4% to 81.2%).
@@
To address this issue, we introduce Fine-grained CLIPScore (F-CLIPScore), a novel image-text correlation metric. F-CLIPScore leverages a sentence parser like spaCy spacy and the forward pass of a Vision-Language Model (VLM) like CLIP clip, offering an efficient way to evaluate the vision encoder’s representational capacity. Our experimental results show that applying F-CLIPScore to the OHD-Caps test set improves accuracy by +39.6% without additional training. This indicates that the limited capacity of the vision encoder may not be the primary cause of object hallucination. Additionally, we verify that using F-CLIPScore for pretraining data curation in LVLMs enables the training of models with reduced hallucination, even with significantly fewer data samples. Notably, in LVLM pretraining, data filtering alone improved POPE accuracy by 4.9% compared to the baseline.
@@
  • We provide evidence suggesting that, in discriminative evaluation, object hallucination is not primarily explained by the vision encoder’s representational capacity alone.
@@ 2 Related Work
Object hallucination refers to cases in which the generated textual descriptions include objects that do not correspond to the given image (lvlm_hallucination_survey). LVLMs generally consist of three components: a vision encoder, an LLM, and an adapter (LLaVA). The structural characteristics of LVLMs contribute to object hallucination, which arises from multiple intertwined factors (lvlm_hallucination_survey). While some studies argue that hallucinations can be mitigated by enhancing the decoding process of the LLM (lcd_lvlm; icd_lvlm), others suggest that one of the causes lies in the limited representational capacity of the vision encoder (ohd-caps). Additionally, some research indicates that training the adapter with contrastive data is essential to reduce object hallucination (hacl). Moreover, the trained bias of the model has also been identified as a cause of hallucination (ciem; robust_instruction_tuning).
CLIPScore (clipscore) is a reference-free evaluation metric that assesses the consistency between an image and text caption by computing the cosine similarity between the embeddings generated by the vision encoder and text encoder of the CLIP model. Beyond its application in measuring caption quality, several studies have also leveraged CLIPScore for data curation in the training of Vision-Language Models (VLMs) (laion; datacomp).
A recent study utilized CLIPScore to evaluate object hallucination in Vision-Language Models (VLMs) (ohd-caps). Their findings suggest that this phenomenon stems from the limited capacity of the vision encoder. In this study, we carefully reassess this claim and demonstrate that object hallucination is not necessarily caused by the limitations of the vision encoder alone.
@@ 3.1 Motivation
We first computed the distribution of cosine similarity between the image embeddings from CLIP-L and OHD-Caps-trained CLIP-L, focusing on the samples that were correctly predicted by CLIP-L. We then performed the same analysis on the samples that were incorrectly predicted by CLIP-L and compared the two distributions. As shown in Figure 2a, we observe that there is no substantial difference between the two distributions. We conducted the same analysis using ground-truth text embeddings of the pre-trained and fine-tuned models, and similarly found no significant difference in the cosine similarity distributions between the samples that were correctly and incorrectly predicted by CLIP-L (Figure 2b). Two-sample t-tests yield p-values of $0.11$ for a and $0.67$ for b. In contrast, Figure 2c reveals significant changes in text embeddings for captions with hallucination (green), and without hallucination (purple) highlighting the distinct adaptation of text representation. Two-sample t-tests yield near zero p-value for c. These observations suggest that OHD-Caps training induces more noticeable changes in the text representation space, particularly in discriminating hallucinated from non-hallucinated captions, whereas the vision representation space exhibits relatively minor changes.
@@ 3.2 Fine-grained CLIPScore
Motivated by these observations, we propose a simple metric called Fine-grained CLIPScore (F-CLIPScore), which enhances the discriminative power of the VLM in order to utilize textual information with more granularity without additional training. F-CLIPScore first utilizes the spaCy parser spacy to extract nouns from a given sentence. Then, it evaluates the quality of an image caption by averaging the CLIPScore of the entire sentence and each individual noun (Figure 3).
Mathematically, given an image $I$ and a caption $C$ containing a total of $N$ nouns, each denoted as $n_{i}$, F-CLIPScore is defined as Eq. 1.
$\small\operatorname{F-CLIPScore}(I,C)=\frac{\operatorname{CLIPScore}(I,C)+\sum_{i=1}^{N}{\operatorname{CLIPScore}(I,n_{i})}}{N+1}$ (1)
@@ 4 Experiments
For evaluating OHD-Caps test set, F-CLIPScore only requires an additional parsing step (we used en_core_web_sm of spaCy (spacy)), which takes on average 170 ms per sentence and can be parallelized across samples. We note that this parser runs in linear time with respect to the length of the sentence (spacyparser). For training CLIP with Eq. 2 on the OHD-Caps train set, we used a batch size of 64 and a learning rate of 1e-5, which required 3 hours on an H100 GPU. For LLaVA pretraining, we employed an effective batch size of 256 and a learning rate of 1e-3, which took 7 hours on two H100 GPUs.
@@ 5.1 F-CLIPScore on OHD-Caps
We evaluated the OHD-Caps test set, an object hallucination assessment dataset, using the proposed F-CLIPScore. As shown in Table 1, evaluation results with OpenAI CLIP ViT-L clip indicate that F-CLIPScore outperformed the baseline model by up to 39.6% without additional training (row 4 vs. 5). However, it still performed 17.6% to 37.4% worse than trained models (row 5 vs. 7). This may indicate that although CLIP vision encoders may not be the main cause of object hallucination, further training may enhance their capability.
@@
$\dagger$: We trained openai/clip-vit-large-patch14 with the train code from ohd-caps
@@
To test whether F-CLIPScore is orthogonal to the previously proposed method ohd-caps, we modify the loss as
@@
where $L_{CLIP}$ is the contrastive loss proposed in  clip, and $L_{OHD}$ is the marginal loss proposed in  ohd-caps. $B$ denotes the batch size, while $I_{i}$ and $C_{i}$ are image and caption of the $i$-th positive pair. $\alpha$ is a hyperparameter that we set to $0.3$. Applying F-CLIPScore as a loss improved performance by 0.7 percentage points on COCO and 0.1 percentage points on NoCaps (row 7 vs. 9). These results suggest that F-CLIPScore could be a complementary component in VLM training.
We randomly replaced the caption embeddings into noun embeddings extracted from a Wikipedia corpus (davies2015wikipedia) to examine how F-CLIPScore responds to such perturbations. As shown in Table 2, we observed the expected decline in accuracy as nouns were replaced. This indicates that the strong performance of F-CLIPScore, even without additional training, is not due to random noise, but rather stems from its ability to more effectively leverage fine-grained textual embeddings.
@@
Table 3: Accuracy on the OHD-Caps test set evaluated with OpenAI CLIP-L. $\operatorname{F-CLIPScore}_{N}$ refers to the noun-based $\operatorname{F-CLIPScore}$ defined in Eq. 1. $\operatorname{F-CLIPScore_{NP}}$ uses noun phrases instead of individual nouns, and $\operatorname{F-CLIPScore}_{V}$ replaces the nouns with verbs.
@@
CECLIP (ceclip) 32.8 28.0 25.0
NegCLIP (negclip) 52.8 40.8 23.4
@@
Our F-CLIPScore even outperforms the training-based methods (ceclip; negclip), as shown in Table 4.
@@
SigLIP ViT-L (CLIPScore) (siglip) 48.6 38.4 30.2
@@
5.2 LVLM Pretrain Data Curation with F-CLIPScore
As shown in Section 5.1, F-CLIPScore was able to exhibit competent performance in detecting object hallucination without any training, and could further serve as an complementary method for training VLMs. We aimed to investigate whether F-CLIPScore can influence the pretraining process of LVLM by acting as a data filter between the vision encoder and the LLM. To this end, we applied F-CLIPScore to filter the pretraining data for LLaVA (LLaVA), which consists of 558k samples, by removing the bottom $x\%$ of the alignment training set based on F-CLIPScore.
POPE Acc ($\uparrow$, %) Filtering rate (%)
Filtering Method 20 30 40 50 60
Random 50.7 49.9 49.9 50.5 49.8
CLIPScore (w/o train) 52.2 47.3 50.0 53.1 50.0
F-CLIPScore (w/o train) 51.6 55.5 50.6 50.1 51.8
CLIPScore (trained) 49.8 49.8 50.8 49.8 52.6
w/o filtering 50.6
Table 6: Accuracy on the POPE benchmark after LLaVA pretraining (vision-text alignment training before SFT). We use CLIP-L clip as a vision encoder, and Llama 2 7B llama2 as an LLM backbone. “trained” indicates the OHD-Caps-trained CLIP-L. “Random” refers to the removal of $x\%$ of samples chosen at random.
As shown in Table 6, training the alignment model on the top 70% of data curated by F-CLIPScore resulted in a +4.9% improvement in POPE accuracy compared to training on the entire dataset (row 5 vs. 7). In contrast, using the OHD-Caps-trained CLIP for filtering yielded only marginal gains, and random filtering showed no improvement. These results may suggest that F-CLIPScore effectively captures object hallucination-related quality, even in general-purpose datasets. This finding underscores the need to explore alternative causes of object hallucination beyond the capacity of the vision encoder. Although model performance shows a non-linear trend with respect to the filtering ratio, our experiments used a fixed number of training epochs, which may account for this behavior. As Goyal_2024_CVPR highlights the need to balance data quality with computational cost, identifying the optimal filtering ratio under such trade-offs remains an important direction for future research.
6 Conclusion
Limitations
First, our study primarily targets discriminative evaluation and retrieval-style scoring (e.g., OHD-Caps, POPE) and does not directly evaluate hallucination in free-form caption generation. Investigating how noun-level scoring can be integrated into generative decoding or training objectives remains future work.
Second, our method faces linguistic constraints and challenges in multilingual generalization. This study employs the spaCy parser (spacy) to extract nouns from text, a technique that performs relatively reliably in well-structured languages such as English. However, parsing accuracy may vary across different languages, potentially leading to inconsistencies in F-CLIPScore computation. To address this, future research should explore the scalability of F-CLIPScore by evaluating its effectiveness on multilingual datasets and refining the parsing methodology for broader linguistic applicability.
Acknowledgments
This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government(MSIT) (RS-2025-23524855).
Appendix A Qualitative Analysis of F-CLIPScore Filtering
We first observed the degree of overlap between the F-CLIPScore and CLIPScore metrics when filtering the data. At filtering rates of 20, 30, 40, 50, and 60, the proportions of overlapping image–caption pairs filtered by both metrics were 57%, 62%, 67%, 71%, and 76%, respectively. We then randomly sampled 10 image–caption pairs that did not overlap between the two metrics’ filtered sets at the 30% filtering rate, which was the setting that yielded the best performance when using the F-CLIPScore metric. The results are presented in Figure 4. As shown in Figure 4, the images filtered by CLIPScore (upper row) appear to include some that should not have been filtered out, despite having normal and semantically correct captions. Although cases such as (g) and (h) involve repetitive lexical usage—where a word is repeated even if it correctly refers to an object present in the image—it is still reasonable for such captions to receive lower scores. However, it is unfortunate that these samples were filtered only by CLIPScore and not by F-CLIPScore.
On the other hand, the images filtered by F-CLIPScore (lower row in Figure 4) include cases such as (l), where even though there is repetitive lexical usage, the corresponding object in the image is actually a model rather than a designer, making the filtering reasonable. There were also examples like (n), where the captioned object is missing from the image, and (r), where the image shows a personal debt chart for “Brainy” rather than a “brain”. These qualitative samples provide insights into how filtering with F-CLIPScore influenced the performance of the LVLMs.
(a) a view of the crowd during the 2016 nba allstar game at madison arena on august 30 2018 in new
(b) a group of food vectors on a white background
(c) a pair of samsung charger stand with charging and charger in the background
(d) an illustration of lord krishna in marathi language
(e) the duke and duchess of cambridge pose for a family portrait
(f) blue and yellow macaw in the jungle sticker
(g) a cliff cliff is on the island of maui canvas bag
(h) the brown shoe is shown with two black soles and a black rubber outsole
(i) garmin edge 20 gps watch with heart rate and gps coordinates
(j) ayrmoors - long length coat 8
(I) Image samples filtered by CLIPScore at the 30% filtering rate
(k) dermastir’d® lux ultra k wax cream spf
(l) an elegant black designer designer gown with beautiful lace work
(m) pdf manual for microphone microphones for alto telephone
(n) a stainless steel sink with three drainers and two spout faucets
(o) horizon™ magico mesh coil
(p) tv on the stone beach on sunset stock photo
(q) battlestard galactica the complete series blu - ray
(r) an upward chart shows the number of complaints about brain disease
(s) the official documents of the commission for commissioning and remediarnation of power transmission units.png
(t) telefish garden pitcher in new castle pa, the bouquet
(II) Image samples filtered by F-CLIPScore at the 30% filtering rate
Figure 4: Randomly sampled 10 image-caption pairs each from (I) samples filtered by CLIPScore and (II) samples filtered by F-CLIPScore at the 30% filtering rate. Images overlapping between the two metrics were excluded.
Furthermore, we measured the CLIPScore and F-CLIPScore for each image–caption pair across the entire LLaVA-Pretrain dataset and sorted them in ascending order, such that a higher score corresponds to a higher rank. We then sorted the samples by the difference between the two ranks (F-CLIPScore rank - CLIPScore rank) and selected the top 10 samples with the largest positive values (i.e., those that CLIPScore rated higher than F-CLIPScore, shown in the upper row of Figure 5) and the bottom 10 samples with the largest negative values (i.e., those that F-CLIPScore rated higher than CLIPScore, shown in the lower row of Figure 5).
As shown in Figure 5, the samples on side (I), which CLIPScore rated higher than F-CLIPScore, contain many text-rendered images. Conversely, the samples on side (II), which F-CLIPScore rated higher than CLIPScore, tend to exhibit duplicated lexical usage. These findings are consistent with the previous observations, and such a trade-off is expected to make determining an appropriate filtering rate more challenging.
(a) the image shows a photo of a money trap with text escaping the debt trap mondaysmoneyuk twitter chat
(b) an elderly couple with a check presentation from the penns valley ambulance in michigan
(c) a lake michigan skyline and a chicago skyline with the text the ultimate list of top paid summer internship opportunities in
(d) nikwax tech wash for bike helmets, helmets and helmets
(e) a knife, a clochet, and a round-up of the nobox household appliances gift give
(f) samsung a70 battery drain test, testing and test battery at home in 4 days | best battery indicator
(g) a girl lying on a skateboard with the words holiday modern mobile game guide in front
(h) two men shaking hands outside of a plane with the title, 5 ways to market your private jet charter
(i) a laptop, phone, coffee cup, and keyboard with the words 5 things every community bank needs
(j) a calendar with the words legacy family tree webinars for a holiday meal
(I) Top 10 samples with the largest positive ranking differences
(k) solitaire solita solitaire solita solitaire solitaire solitaire solitaire solitaire solitaire sol
(l) alexander alexander - double breasted coat 9
(m) a screenplay screenplay with scriptwriting, handwritten
(n) an illusion illusion illusion illusion in black and white
(o) a52 250gb sat - 540 internal ssd
(p) a fujifilm m7 with the fuji 105mm f2
(q) the gap and the gap of the gaps gap of the gaps gap on the gaps gap of the gaps is
(r) the headstone of robert s dutton
(s) old cassette tape recorder
(t) the algorithm, with an allusion
(II) Bottom 10 samples with the largest negative ranking differences
Figure 5: The figure shows the top and bottom 10 samples from the entire LLaVA-Pretrain dataset, sorted by the difference between the F-CLIPScore rank and the CLIPScore rank. (I) represents samples that CLIPScore rated higher than F-CLIPScore, while (II) represents the opposite cases. Each caption is written below its corresponding image.
Appendix B Licenses for the datasets and models used